OpenStudy (anonymous):

The number of traffic violations in a city increased from 12 per month to 16 per month over the course of two years. What is the percent of increase in the number of traffic violations per month for that two-year period? Round to the nearest percent, if necessary.

OpenStudy (rulnick):

(new - old) / old = (16-12)/12 = 4/12 = 1/3 = 33 and 1/3 percent